OpenStudy (anonymous):

Any tips for foiling numbers with a coefficient in front of the x^2? (like 2x^2 + x -3). I can only get them with "guess and check"

OpenStudy (anonymous):

multiply the numbered coefficient across . (3x^2-5)(x^2+6) =3x^4+18x^2-5x^2-30

OpenStudy (anonymous):

I know how to factor them out of parenthesis... I just don't really know how to get them back in, if that makes any sense.

OpenStudy (anonymous):

Like, I know that (3x^2-2)(x+1)= 3x^2-x-2, but I don't know how to put 3x^2 -x -2 into parenthesis... the 3 messes everything up.

OpenStudy (anonymous):

oh okay. well multiply the whole thing back in i.e. 4x^2(2x-5) 8x^3-20x^2 variables multiply with variables and numbers multiply with numbers also when multiplying powers such as x^2 and x^3 you add the exponent =x^5

OpenStudy (anonymous):

wrote that wrong... the 3x shouldn't have been squared

OpenStudy (anonymous):

in the parenthesis.

OpenStudy (anonymous):

let me restate... say that you have a simple equation, like x^2+ 2x +1. I know you look at all possible pairs of multiples that equal 1, and then see if any can add or subtract to equal two. But if it's 4x^2 + 2x+ 1, then the add/ subtract multiples of 1 don't matter because the 4 messes things up

OpenStudy (anonymous):

the best way to factor out equations is to just plug and chug. trial and error worked best for me. do it enough and youll get fast at it. there may be ways to do it quicker but sometimes the more you try to split it up the nastier it gets.

OpenStudy (anonymous):

oh. Thanks. Sometimes trial and error really is the best way- basically what evolution is right?

OpenStudy (anonymous):

yeah. hopefully your teach wont be a douche and will make it fairly obvious when stating the problem. anyways good luck